OpenStudy (lisa123):

Factor x^3-4x^2+5x-2

MsBrains (ms-brains):

This might help

OpenStudy (anonymous):

help me please

OpenStudy (lisa123):

how did you get (x^2-3x+2)(x-1)?

OpenStudy (jchick):

OOps wrong question lol I will help with yours

OpenStudy (jchick):

x^3 – 4x^2 + 5x – 2

OpenStudy (jchick):

That is what you have right?

OpenStudy (lisa123):

yes

OpenStudy (jchick):

ok

OpenStudy (jchick):

Note if (x + a) is a factor, then a will be a factor of -2. Those are ±1 or ±2.

OpenStudy (jchick):

Note that if x = 1, then 1^3 – 4(1)^2 + 5(1) – 2 = 1 – 4 + 5 – 2 = 0, so x – 1 is a facto

OpenStudy (lisa123):

ok

OpenStudy (jchick):

factor.*

OpenStudy (jchick):

Therefore x^3 – 4x^2 + 5x – 2 = (x – 1)(x² – 3x + 2)

OpenStudy (jchick):

Now I recognize that (x² – 3x + 2) = (x – 1)(x – 2)

OpenStudy (jchick):

x^3 – 4x^2 + 5x – 2 = (x – 1)²(x – 2)

OpenStudy (jchick):

Is there a specific law you were supposed to use with this?

OpenStudy (lisa123):

I think its the rational roots thm

OpenStudy (jchick):

Ok

OpenStudy (lisa123):

thank you! @jchick

OpenStudy (jchick):

No problem!

OpenStudy (lisa123):

got it!

OpenStudy (jchick):

Sorry typo (x-2) (x-1)^2

OpenStudy (jchick):

There is your final answer!